OpenStudy (anonymous):

Work-Out Corner has 5 more than 3 times as many exercise bicycles as The Gym. Together they have 21 bicycles. Solve the equation x + 3x +5 = 21 to find the number of bicycles at Work-Out Corner. (1 point) 4 bicycles 17 bicycles 7 bicycles 25 bicycles

OpenStudy (anonymous):

(workout corner)-5=3*(the gym)

OpenStudy (anonymous):

4x+5=21 4x=16 x=4

OpenStudy (anonymous):

Thank you! :3